OpenStudy (anonymous):

How do you solve sin^-1 ((√2)/2)?

OpenStudy (saifoo.khan):

Can you use a calculator?

OpenStudy (anonymous):

\[\sin^{-1} (\frac{ \sqrt{2} }{ 2 })\]

OpenStudy (ash2326):

@Insist do you know the value of sin 45 degrees?

OpenStudy (anonymous):

A simple one

OpenStudy (anonymous):

How do you find it. o;

OpenStudy (goformit100):

Trigonometry is so easy :)

OpenStudy (ash2326):

@Insist \[\sin 45=\frac 1{\sqrt 2}\] Could you simplify this? \[\frac{\sqrt 2}2\]

OpenStudy (anonymous):

You can't simplify that, but how did you find sin 45

OpenStudy (ash2326):

\[\frac{\sqrt 2}{2}=\frac 1 {\sqrt 2}\]

OpenStudy (ash2326):

sin 45, it's a standard value of trigonometric ratios. It's the sin for a right triangle with equal sides